Determine if the matrix p=
D= 1/√2, 0, 1/√2
0, 1, 0
-1/√2, 2, 1/√2
Is othognal.
1
Expert's answer
2019-09-20T09:16:12-0400

This matrix would be orthogonal if a product of the matrix and its transpose is a unit matrix:


DDT = E


Thus, we should first found a transpose matrix:


"D=
\begin{vmatrix}
\frac{1}{\sqrt{2}} & 0 & \frac{1}{\sqrt{2}} \\
0 & 1 & 0 \\
\frac{-1}{\sqrt{2}} & 2 & \frac{1}{\sqrt{2}}
\end{vmatrix}
D^T=
\begin{vmatrix}
\frac{1}{\sqrt{2}} & 0 & \frac{-1}{\sqrt{2}} \\
0 & 1 & 2 \\
\frac{1}{\sqrt{2}} & 0 & \frac{1}{\sqrt{2}}
\end{vmatrix}"


Now, we must multiply them together:


"DD^T=
\begin{vmatrix}
\frac{1}{\sqrt{2}} & 0 & \frac{1}{\sqrt{2}} \\
0 & 1 & 0 \\
\frac{-1}{\sqrt{2}} & 2 & \frac{1}{\sqrt{2}}
\end{vmatrix}
\begin{vmatrix}
\frac{1}{\sqrt{2}} & 0 & \frac{-1}{\sqrt{2}} \\
0 & 1 & 2 \\
\frac{1}{\sqrt{2}} & 0 & \frac{1}{\sqrt{2}}
\end{vmatrix}
=
\begin{vmatrix}
1 & 0 & 0 \\
0 & 1 & 2 \\
0 & 2 & 5
\end{vmatrix}"


The product is not the unit matrix. Thus, the D matrix was not orthogonal.